# Bayesian Inference for Relational Graph in a Causal Vector Autoregressive Time Series

### v1

@@ 3.1 Efficient Computation of the Likelihood
The computation of the likelihood is difficult for a multivariate time series. For a Gaussian VAR process, the likelihood can be computed relatively fast using the Durbin-Levison (DL) or innovations algorithm. However, under different parameterizations, the computational burden can increase significantly depending on the complexity of the parameterization. Moreover, the DL-type algorithm can still be challenging if the process dimension is high.
In models with a high number of parameters, a common approach is to seek low-dimensional sub-models that could adequately describe the data and answer basic inferential questions of interest.
Using the proposed parameterization of $\bm{\Omega}$ and the differences $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$, we provide a low-rank formulation of a causal VAR process that leads to an efficient algorithm for computing the likelihood based on a Gaussian VAR($p$) sample. The algorithm achieves computational efficiency by avoiding inversion of large dimensional matrices. The sparse precision matrix $\bm{\Omega}=\bm{C}_{0}^{-1}$ is modeled as a separate parameter, allowing direct inference about the graphical structure under sparsity constraints . The likelihood computation requires the inversion of only $r_{j}\times r_{j}$ matrices instead of $d\times d$ matrices. The last fact substantially decreases the computational complexity. In particular, when $r_{j}=r=1$, i.e. when the conditional precision updates, $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$, are all rank-one, all components of likelihood computation can be done without matrix inversion or factorization.
@@
Specifically, for each $j=1,\ldots,p,$ along with $\bm{L}_{j}$ we will use $d\times r_{j}$ parameters arranged in a $d\times r_{j}$ matrix, $\bm{K}_{j}$, to describe the reduced-rank updates $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$. The exact definition of the matrices $\bm{K}_{j}$ along with the steps for computing the causal VAR likelihood based on the basic constraint-free parameters $\bm{\Omega}$, $\bm{L}_{j}$, and $\bm{K}_{j}$, $j=1,\ldots,p$ are given below. We assume the that the ranks $r_{1},\ldots,r_{p}$ are specified and fixed. Also, unless otherwise specified, we will assume the matrix square roots for pd matrices to be the unique symmetric square-root. Then, given the parameters $\bm{\Omega},\bm{L}_{1},\ldots,\bm{L}_{p}$ and the initialization $\bm{C}_{0}^{-1}=\bm{\Omega},$ the following steps describe the remaining parameters $\bm{K}_{1},\ldots,\bm{K}_{p}$ recursively along with recursive computation of the components $f(\bm{X}_{j}|\bm{Y}_{j-1})$ of the VAR likelihood.
@@
3.2 Identification of the Parameters
In the recursive algorithm described above, the parameters $\bm{\Omega},(\bm{L}_{1},\bm{K}_{1}),\ldots,(\bm{L}_{p},\bm{K}_{p})$ are mapped to the modified parameters $\bm{\Omega},(\bm{U}_{1},\bm{V}_{1}),\ldots,(\bm{U}_{p},\bm{V}_{p})$. The mapping is not one-to-one since each $\bm{V}_{j}$ has the restriction $\bm{V}_{j}^{\mathrm{T}}\bm{D}_{j-1}^{-1}\bm{V}_{j}=\bm{I}$ while the basic parameters $\bm{K}_{j}$ do not have any restrictions. The intermediate parameters in the mapping, $\bm{\Omega},(\bm{U}_{1},\bm{V}_{1}),\ldots,(\bm{U}_{p},\bm{V}_{p})$, are identifiable only up to rotations. Consider the equivalence classes of pair of $d\times r_{j}$ matrices $(\bm{U},\bm{V})$ defined through the relation $(\bm{U},\bm{V})\equiv(\bm{U}\bm{Q},\bm{V}\bm{Q})$ for any $r_{j}\times r_{j}$ orthogonal matrix $\bm{Q}$. Let for each $r_{j}$, $\mathbb{C}_{r_{j}}$ be the set of such equivalence classes. Also let $\mathcal{S}^{++}_{d,s}$ be a subset of the positive definite cone $\mathcal{S}^{++}_{d}$ with specified sparsity level $s$. The following proposition shows that there is a bijection from the set $\mathcal{S}^{++}_{d,s}\times\mathbb{C}_{r_{1}}\times\cdots\times\mathbb{C}_{r_{p}}$ to the causal VAR$(p)$ parameter space defined by the VAR coefficients $\bm{A}_{1},\ldots,\bm{A}_{p}$ and innovation variance $\bm{\Sigma}$, where polynomials with coefficients $\bm{A}_{1},\ldots,\bm{A}_{p}$ satisfy low-rank restrictions and the stationary precision satisfies $s$-sparsity. The definition of sparsity is made more clear in the prior specification section. For simplicity, we use a mean zero VAR($p$) process. Before we formally state the identification result, we state and prove a short Lemma that is used to show identification.
Lemma 1.
Given a $n\times n$ matrix $\bm{A}$ of rank $1\leq r\leq n$ for some positive integer $n$, and a positive definite $n\times n$ matrix $\bm{B}$, there exists $n\times r$ matrices $\bm{C}$ and $\bm{R}$ such that $\bm{A}=\bm{C}\bm{R}^{\mathrm{T}}$, $rank(\bm{C})=rank(\bm{R})=r$ and $\bm{R}^{\mathrm{T}}\bm{B}\bm{R}=\bm{I}$
Proof.
Given a rank factorization $\bm{A}=\bm{U}\bm{V}$ and the symmetric square root $\bm{B}^{1/2}$, let $\bm{S}=\bm{Q}\bm{Z}$ be the Q-R decomposition of $\bm{S}=\bm{B}^{1/2}\bm{V}^{\mathrm{T}}$ where $\bm{Q}$ is $n\times r$ semi-orthogonal and $\bm{Z}$ is $r\times r$ nonsingular matrix. Then $\bm{A}=\bm{U}\bm{V}=\bm{U}\bm{V}\bm{B}^{1/2}\bm{B}^{-1/2}=\bm{U}\bm{V}\bm{B}^{1/2}\bm{B}^{-1/2}=\bm{U}\bm{Z}^{\mathrm{T}}\bm{Q}^{\mathrm{T}}\bm{B}^{-1/2}=\bm{C}\bm{R}$ where $\bm{C}=\bm{U}\bm{Z}^{\mathrm{T}}$ and $\bm{R}=\bm{B}^{-1/2}\bm{Q}$. For this choice, $\bm{R}^{\mathrm{T}}\bm{B}\bm{R}=\bm{I}.$
Proposition 2.
Let $\bm{\Omega}>0$ be a $d\times d$ positive definite matrix, which is $s$-sparse. Let $(\bm{U}_{1},\bm{V}_{1}),\ldots$, $(\bm{U}_{p},\bm{V}_{p})$ be given full column rank matrix pairs, of order $d\times r_{1},\dots,d\times r_{p}$, respectively. Then there is a unique causal $\mathrm{VAR}(p)$ process with stationary variance matrix $\bm{\Gamma}(0)=\bm{\Omega}^{-1}$ and autcovariances $\bm{\Gamma}(1),\ldots,\bm{\Gamma}(p)$ (and hence the coefficients $\bm{A}_{1},\ldots,\bm{A}_{p}$ as the Yule-Walker solution) uniquely determined recursively from $\bm{W}_{j}=\bm{U}_{j}\bm{V}_{j}^{\mathrm{T}}$. The associated increments in the conditional precision matrices $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$ will be of rank $r_{j}.$
Proof.
The over-parameterization of the reduced rank matrices is intentional; it reduces computational complexity. Consistent posterior inference on the set of identifiable parameters of interest is still possible, and the details are given in subsequent sections. For a general $d\times d$ matrix with rank $r$, the number of free parameters is $d^{2}-(d-r)^{2}=2dr-r^{2}$ where $(d-r)^{2}$ is the co-dimension of the rank manifold of $d\times d$ rank $r$ matrices. In the proposed parameterization, we are parameterizing the rank $r_{j}$ updates $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$ using $2dr_{j}$ parameters in the matrix pair $(\bm{L}_{j},\bm{K}_{j}).$ Thus, we have $r_{j}^{2}$ extra parameters that are being introduced for computational convenience. Typically, $r_{j}$ will be small and hence the number of extra parameters will be small as well.
@@ 3.3 Rank-one Updates
The case when the increments $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$ are rank one matrices, i.e. $r_{j}=1$ for all $j$, is of special interest. Then, the parameters $\bm{U}_{j},\bm{V}_{j}$ are fully identifiable if one fixes the sign of the first entry in $\bm{U}_{j}$. Moreover, in this case, the likelihood can be computed without having to invert any matrices or computing square roots.
@@ 3.4 Prior Specification
If the model is assumed to be causal, the prior should charge only autoregression coefficients complying with causality restrictions while putting prior probability directly on the stationary covariance matrix. Presently, the literature lacks a prior that charges only causal processes with the prior on the precision $\bm{\Omega}$ as an independent component of the prior. Using our parametrizations, we achieve that. Recall that in our setup, $p$ is considered given; hence, no prior distribution is assigned to $p$. Also, the ranks $r_{1},\ldots,r_{p}$ are specified. While all of the procedures described in this paper go through where the $p$ ranks are potentially different from each other, we develop the methodology of low-rank updates by fixing $r_{1}=\cdots=r_{p}=r.$
We consider the modified Cholesky decomposition of $\bm{\Omega}=(\bm{I}_{p}-\bm{E})\bm{F}(\bm{I}_{p}-\bm{E})^{\mathrm{T}}$, where $\bm{E}$ is strictly lower triangular, $\bm{F}=\hbox{diag}(\bm{f})$ is diagonal with $\bm{f}$ the vector of entries. For sparse estimation of $\bm{\Sigma}^{-1}$, we put sparsity inducing prior on $\bm{E}$. We first define a hard thresholding operator $H_{\lambda}(\bm{H})=\{h_{i,j}\mathbbm{1}(|h_{i,j}|>\lambda)\}$. Then we set, $\bm{E}=H_{\lambda}(\bm{E}_{1})$, where $\bm{E}_{1}$ is a strictly lower triangular matrix. For off-diagonal entries $\bm{E}_{1}$, we let $e_{ij,1}\sim\mathrm{N}(0,\sigma^{2}_{e})$ and $\sigma^{2}_{e}\sim\hbox{Inv-Ga}(c_{1},c_{1})$ for $i<j$. The components $f_{1},\ldots,f_{p}$ of $\bm{f}$ are independently distributed according to the inverse Gaussian distributions with density function $\pi_{d}(t)\propto t^{-3/2}e^{-(t-\xi)^{2}/(2t)}$, $t>0$, for some $\xi>0$; see (Chhikara, 1988). This prior has an exponential-like tail near both zero and infinity. We put a weakly informative mean-zero normal prior with a large variance on $\xi$. While employing sparsity using hard thresholding, the modified Cholesky form is easier to work with. Lastly, we put a Uniform prior on $\lambda$.
@@
    where $\lceil x\rceil$ stands for the ceiling function that maps to the smallest integer, greater than or equal to $x$, and Ga stands for the gamma distribution. The parameters $\phi_{lk}$ control local shrinkage of the elements in $\Lambda$, whereas $\tau_{k}$ controls column shrinkage of the $k$-$th$ column. Similarly, $\psi_{k,\lceil{\ell}/{r}\rceil}^{-1})$ helps to shrink higher-order columns in $\bm{L}_{k}$. Following the well-known shrinkage properties of multiplicative gamma prior, we let $\kappa_{1}=2.1$ and $\kappa_{2}=3.1$, which work well in all of our simulation experiments. However, in the case of rank-one updates, the $\psi_{k,m}$’s are omitted. Computationally, it seems reasonable to keep $\psi_{k,m}=1$ as long as $r$ is pre-specified to a small number.
  • $\bm{K}_{j}$ of dimension $d\times r$: We consider non-informative flat prior for the entries in $\bm{K}_{j}$. The rank of $\bm{K}_{j}$ is specified simultaneously with the rank of $\bm{L}_{j}$. Hence, the cumulative shrinkage prior for $\bm{K}_{j}$ is not needed.
@@ 3.5 Posterior Computation
Adaptive Metropolis-Hastings (M-H) moves are used to update the lower-triangular entries in the latent $\bm{E}_{1}$ and the entries in $\bm{f}$. Due to the positivity constraint on $\bm{f}$, we update this parameter in the log scale with the necessary Jacobian adjustment. Specifically, we generate the updates from a multivariate normal, where the associated covariance matrix is computed based on the generated posterior samples. Our algorithm is similar to Haario et al. (2001) with some modifications, as discussed below. The initial part of the chain relies on random-walk updates, as no information is available to compute the covariance. However, after the 3500th iteration, we start computing the covariance matrices based on the last $S$ accepted samples. The choice of 3500 is based on our extensive simulation experimentation. Instead of updating these matrices at each iteration, we update them once at the end of each 100-$th$ iteration using the last $S$ accepted samples. The value of $S$ is increased gradually. Furthermore, the constant variance, multiplied with the covariance matrix as in Haario et al. (2001) is tuned to maintain a pre-specified level of acceptance. The thresholding parameter $\lambda$ is also updated in the log scale with a Jacobian adjustment.
@@
Recently, Heaps (2023) applied the parametrization from Roy et al. (2019) with flat priors on the new set of parameters and developed HMC-based MCMC computation using rstan. Our priors, however, involve several layers of sparsity structures and, hence, are inherently more complex. We primarily rely on M-H sampling, as direct computation of the gradients that are required for HMC is difficult. We leave the exploration of more efficient implementation using rstan to implement gradient-based samplers as part of possible future investigations.
@@ 4 Posterior Contraction
To prove the theorem, we apply Theorem 3 of Ghosal and Van Der Vaart (2007) (equivalently, Theorem 8.19 of Ghosal and van der Vaart (2017)) and verify the testing condition directly. As the joint distribution of all components of the observations is a $dT$-dimensional multivariate normal, likelihood ratios can be explicitly written down. We construct a test based on the likelihood ratio at a few selected alternative values and obtain uniform bounds for likelihood in small neighborhoods of these alternative values to establish that the resulting test also has exponentially low type II error probabilities in these small neighborhoods. The resulting finitely many tests are combined to obtain the desired test, a technique used earlier in the high-dimensional context by Ning et al. (2020); Jeong and Ghosal (2021); Shi et al. (2021) for independent data. By relating the Kullback-Leibler divergence for the time series to the Frobenius distance on the precision matrix $\bm{\Omega}$ and the low-rank increment terms $\bm{L}_{1},\ldots,\bm{L}_{p},\bm{K}_{1},\ldots,\bm{K}_{p}$, we obtain the prior concentration rate, as in Lemma 8 of Ghosal and Van Der Vaart (2007) for a one-dimensional Gaussian time series. The details are shown in the supplementary material section.
@@
5 Numerical Illustrations
We present the results from a limited simulation experiment and also the results from applying the proposed method to analyze graphical structures among several components time series for the US gross domestic product (GDP).
5.1 Simulation

### v2

@@ 4.1 Efficient Computation of the Likelihood
The computation of the likelihood is difficult for a multivariate time series. For a Gaussian VAR process, the likelihood can be computed relatively fast using the Durbin-Levison (DL) or innovations algorithm (Brockwell and Davis, 2009). However, under different parameterizations, the computational burden can increase significantly depending on the complexity of the parameterization. Moreover, the DL-type algorithm can still be challenging if the process dimension is high.
In models with many parameters, a common approach is to seek low-dimensional sub-models that could adequately describe the data and answer basic inferential questions of interest.
Using the proposed parameterization of $\bm{\Omega}$ and the differences $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$, we provide a low-rank formulation of a causal VAR process that leads to an efficient algorithm for computing the likelihood based on a Gaussian VAR($p$) sample. The algorithm achieves computational efficiency by avoiding the inversion of large-dimensional matrices. The sparse precision matrix $\bm{\Omega}=\bm{C}_{0}^{-1}$ is modeled as a separate parameter, allowing direct inference about the graphical structure under sparsity constraints . The likelihood computation requires the inversion of only $r_{j}\times r_{j}$ matrices instead of $d\times d$ matrices. The last fact substantially decreases the computational complexity. In particular, when $r_{j}=r=1$, i.e. when the conditional precision updates, $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$, are all rank-one, all components of likelihood computation can be done without matrix inversion or factorization.
@@
Specifically, for each $j=1,\ldots,p,$ along with $\bm{L}_{j}$ we will use $d\times r_{j}$ parameters arranged in a $d\times r_{j}$ matrix, $\bm{K}_{j}$, to describe the reduced-rank updates $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$. The exact definition of the matrices $\bm{K}_{j}$ along with the steps for computing the causal VAR likelihood based on the basic constraint-free parameters $\bm{\Omega}$, $\bm{L}_{j}$, and $\bm{K}_{j}$, $j=1,\ldots,p$ are given below. We assume the ranks $r_{1},\ldots,r_{p}$ are specified and fixed. Also, unless otherwise specified, we will assume the matrix square roots for pd matrices to be the unique symmetric square-root. Then, given the parameters $\bm{\Omega},\bm{L}_{1},\ldots,\bm{L}_{p}$ and the initialization $\bm{C}_{0}^{-1}=\bm{\Omega},$ the following steps describe the remaining parameters $\bm{K}_{1},\ldots,\bm{K}_{p}$ recursively along with recursive computation of the components $f(\bm{X}_{j}|\bm{Y}_{j-1})$ of the VAR likelihood.
@@
The over-parameterization of the reduced rank matrices is intentional; it reduces computational complexity. Consistent posterior inference on the set of identifiable parameters of interest is still possible, and the details are given in subsequent sections. For a general $d\times d$ matrix with rank $r$, the number of free parameters is $d^{2}-(d-r)^{2}=2dr-r^{2}$ where $(d-r)^{2}$ is the codimension of the rank manifold of $d\times d$ rank $r$ matrices. In the proposed parameterization, we are parameterizing the rank $r_{j}$ updates $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$ using $2dr_{j}$ parameters in the matrix pair $(\bm{L}_{j},\bm{K}_{j}).$ Thus, we have $r_{j}^{2}$ extra parameters that are being introduced for computational convenience. Typically, $r_{j}$ will be small, and hence the number of extra parameters will be small as well.
@@ 4.2 Rank-one Updates
The case when the increments $\bm{C}_{j}^{-1}-\bm{C}_{j-1}^{-1}$ are rank one matrices, i.e. $r_{j}=1$ for all $j$, is of special interest. Then, the parameters $\bm{U}_{j},\bm{V}_{j}$ are fully identifiable if one fixes the sign of the first entry in $\bm{U}_{j}$. Moreover, in this case, the likelihood can be computed without inverting any matrices or computing square roots.
@@ 4.3 Prior Specification
If the model is assumed to be causal, the prior should charge only autoregression coefficients complying with causality restrictions while putting prior probability directly on the stationary covariance matrix. Presently, the literature lacks a prior that charges only causal processes with the prior on the precision $\bm{\Omega}$ as an independent component of the prior. Using our parametrizations, we achieve that. Recall that in our setup, $p$ is considered given; hence, no prior distribution is assigned to $p$. Also, the ranks $r_{1},\ldots,r_{p}$ are specified. While all of the procedures described in this paper go through where the $p$ ranks are potentially different, we develop the methodology of low-rank updates by fixing $r_{1}=\cdots=r_{p}=r$.
We consider the modified Cholesky decomposition of $\bm{\Omega}=(\bm{I}_{p}-\bm{E})\bm{F}(\bm{I}_{p}-\bm{E})^{\mathrm{T}}$, where $\bm{E}$ is strictly lower triangular, $\bm{F}=\hbox{diag}(\bm{f})$ is diagonal with $\bm{f}$ the vector of entries. For sparse estimation of $\bm{\Sigma}^{-1}$, we put sparsity inducing prior on $\bm{E}$. We first define a hard thresholding operator $H_{\lambda}(\bm{H})=(\!(h_{i,j}\mathbbm{1}\{|h_{i,j}|>\lambda\})\!)$. Then we set, $\bm{E}=H_{\lambda}(\bm{E}_{1})$, where $\bm{E}_{1}$ is a strictly lower triangular matrix. For off-diagonal entries $\bm{E}_{1}$, we let $e_{ij,1}\sim\mathrm{N}(0,\sigma^{2}_{e})$ and $\sigma^{2}_{e}\sim\hbox{Inv-Ga}(c_{1},c_{1})$ for $i<j$. The components $f_{1},\ldots,f_{p}$ of $\bm{f}$ are independently distributed according to the inverse Gaussian distributions with density function $\pi_{d}(t)\propto t^{-3/2}e^{-(t-\xi)^{2}/(2t)}$, $t>0$, for some $\xi>0$; see (Chhikara, 1988). This prior has an exponential-like tail near both zero and infinity. We put a weakly informative mean-zero normal prior with a large variance on $\xi$. While employing sparsity using hard thresholding, the modified Cholesky form is easier to work with. Lastly, we put a Uniform prior on $\lambda$.
@@
    where $\lceil x\rceil$ stands for the ceiling function that maps to the smallest integer greater than or equal to $x$, and Ga stands for the gamma distribution. The parameters $\phi_{lk}$ control local shrinkage of the elements in $\Lambda$, whereas $\tau_{k}$ controls column shrinkage of the $k$-$th$ column. Similarly, $\psi_{k,\lceil{\ell}/{r}\rceil}^{-1})$ helps to shrink higher-order columns in $\bm{L}_{k}$. Following the well-known shrinkage properties of multiplicative gamma prior, we let $\kappa_{1}=2.1$ and $\kappa_{2}=3.1$, which work well in all of our simulation experiments. However, in the case of rank-one updates, the $\psi_{k,m}$’s are omitted. Computationally, it seems reasonable to keep $\psi_{k,m}=1$ as long as $r$ is pre-specified to a small number.
  • $\bm{K}_{j}$ of dimension $d\times r$: We consider non-informative flat prior for the entries in $\bm{K}_{j}$. The rank of $\bm{K}_{j}$ is simultaneously specified with that of $\bm{L}_{j}$. Hence, the cumulative shrinkage prior for $\bm{K}_{j}$ is not needed.
@@ 4.4 Posterior Computation
Adaptive Metropolis-Hastings (M-H) moves update the lower-triangular entries in the latent $\bm{E}_{1}$ and the entries in $\bm{f}$. Due to the positivity constraint on $\bm{f}$, we update this parameter in the log scale with the necessary Jacobian adjustment. Specifically, we generate the updates from a multivariate normal, where the associated covariance matrix is computed based on the generated posterior samples. Our algorithm is similar to Haario et al. (2001) with some modifications, as discussed below. The initial part of the chain relies on random-walk updates, as no information is available to compute the covariance. However, after the 3500th iteration, we start computing the covariance matrices based on the last $S$ accepted samples. The choice of 3500 is based on our extensive simulation experimentation. Instead of updating these matrices at each iteration, we update them once at the end of each 100th iteration using the last $S$ accepted samples. The value of $S$ is increased gradually. Furthermore, the constant variance, multiplied by the covariance matrix as in Haario et al. (2001), is tuned to maintain a pre-specified level of acceptance. The thresholding parameter $\lambda$ is also updated on a log scale with a Jacobian adjustment.
@@
Recently, Heaps (2023) applied the parametrization from Roy et al. (2019) with flat priors on the new set of parameters and developed HMC-based MCMC computation using rstan. Our priors, however, involve several layers of sparsity structures and, hence, are inherently more complex. We primarily rely on M-H sampling, as direct computation of the required gradients for HMC is difficult. We leave the exploration of more efficient implementation using rstan to implement gradient-based samplers as part of possible future investigations.
@@ 5 Posterior Contraction
The proofs for Theorem 1 and Theorem 2 are given in the supplementary material section.
6 Differential Connectivity due to the Great Recession Using FRED-QD data
The main objective for developing a flexible stationary graph estimation method for a VAR process is to evaluate how the connectivity between the economic time series in the FRED-QD database changed during the financial crisis of the Great Recession in the context of a stationary graph constructed based on the joint VAR model for the series. The simulation experiment in the next section provides credence to the claim that for graph estimation, there is benefit in explicitly modeling the temporal dependence structure of the sample. As mentioned earlier, a stationary vector autoregressive model is a reasonable model for the FRED-QD time series after performing the transformations suggested in McCracken and Ng (2020). Thus, we estimate a VAR model for the FRED-QD data and the stationary graph obtained from the associated precision matrix.
As argued earlier, fitting two different stationary models to the data from periods before and after the recession will allow us to compare the presumably stationary and potentially different temporal dynamics of the two periods while circumventing the challenging task of accounting for all the changes and associated epochs during the nearly two and half years long period of the Great Recession. Performing a comparison of the stationary graphs between the before and after period in terms of a combined model is relatively straightforward within a Bayesian framework, where we can study the posterior of functions of parameters from the before and after VAR models to evaluate changes in the stationary graph. We specifically study differential connectivity between the FRED-QD series due to the Great Recession in terms of the scale-free edge difference parameters, $\theta_{g}(i,j),$ described in (2.1). We thus fit our proposed model for each group $g$ and each data fragment from before and after the recession. We obtain the posterior samples of stationary precision matrices $\bm{\Omega}_{b,g}$ and $\bm{\Omega}_{a,g}$, before and after the period of great recession, respectively, and construct posterior samples of the scale-free edge difference parameters given in (2.1). As discussed in Section 4.4, we set the maximum VAR order as $\min(p_{m},T/2)$, where $p_{m}=10$ and all the $r_{j}$’s are initialized at 3. Then, at the 1000th iteration, we drop the columns in $\bm{L}_{j}$’s with overall contribution less than 0.1 in the sum of squares. In that process, some of the higher-order $\bm{L}_{j}$’s are dropped, reducing the overall order of the fitted VAR. The computational steps remained stable for higher choices of $p_{m}$ and $r_{j}$’s. All the other tuning details described in Section 4.4 are left unchanged.
To designate the edge-difference parameter as significantly different from zero for any particular pair of economic indicators, and hence determine whether there has been a substantial impact of the great recession on that particular connection, we compute an empirical 95% credible interval for the parameter and check if zero is included in the interval. The particular edge in the stationary graph is inferred to have undergone significant changes during the Great Recession if the credible interval does not contain zero. Let $\{\bm{\Omega}^{(s)}_{b,g}\}_{s=1}^{S}$ be $S$ posterior samples of $\bm{\Omega}_{b,g}$ and similarly $\{\bm{\Omega}^{(s)}_{a,g}\}_{s=1}^{S}$ are posterior samples of $\bm{\Omega}_{a,g}$. For this exercise, we obtain $S=5000$ posterior samples for the before and after precision matrices after burning in the first 5000 samples, thereby obtaining 5000 posterior samples for each $\theta_{g}(i,j)$ parameter. Based on these samples, we compute the 95% credible intervals for $\theta_{g}(i,j)$ as the empirical 2.5 and 97.5 quantiles of the sample of 5000 values. Since there are several hundred pairwise connections, we only report aggregate measures of changes for each of the economic groups in the FRED-QD data. For each group-specific graph, Table 2 reports the percentage of edge-difference parameters with zeroes not included in their credible intervals. Larger values of this number imply a greater disruption of the interdependencies among the series within that group due to the financial crisis. Table 2 values are relatively higher than the ones reported in Table 1, showing that, unlike the analysis based on sparse VAR modeling, the proposed method, based on Bayesian analysis of the difference in the two stationary graphs from before and after periods, detects substantial changes in connectivity within most of the groups. Some of the groups with larger differences in Table 1, such as Stock Markets and Non-Household Balance Sheets, retain higher differences in their stationary precisions in Table 2. However, the inventories, Orders, and Sales series, which showed significantly large differences under he sparse VAR analysis, show only a moderate amount of changes under the proposed detection scheme. Several groups show large differences in their stationary precisions when compared to the estimates in Table 1, e.g., NIPA, Employment, Housing, Earnings and Productivity, Interest Rates, and Money and Credit. The results are consistent with the literature showing substantial changes in the relationship between series from these groups. The groups indicating smaller amount of changes due to the recession under the proposed Bayesian scheme are Industrial Production and Exchange rates. While overall industrial production is significantly affected by the economic downturn, and some relationships are permanently changed, there has been evidence of economic resistance in some parts of the sector that left the interdependencies more comparable between the before and the after recession periods. This is consistent with the literature showing some degree fo resilience and quick recovery in the industrial production series; (Kitsos et al., 2023). A similar picture emerges in the exchange rate group, with only four series. The relatively lower value for the proportion of changes in the exchange rate series may be due to the stability of the safe-haven currencies like the Japanese Yen or the Swiss Franc, or due to the controlled exchange rate dynamics of the Chinese Yuan achieved through stringent financial controls.
To further scrutinize the changes estimated in the graphs for each group, and thereby evaluate the impact of the Great Recession on each of the subgroups, we look at the individual change graph, i.e., a graph of the edges deemed significantly perturbed by the Bayesian method. Figure S4 shows edges $(i,j)$ within each network where the credible intervals of the scale-free edge-difference parameters do not include zero. These edges represent the interdependencies that changed during the great recession.
In each group, several edges are determined to have changed during the Great Recession. This is consistent with the existing knowledge. However, there are series (nodes) for which only a few edges show changes between the before and after recession periods.
In the NIPA group, the Government Consumption, Expenditure, and Gross Investment series (node 13) has only two edges (with export and import series) that show significant change due to the recession. The government consumption and spending worked as a stabilizer during the financial crisis, cushioning the effect of the economic downturn Ramey (2019). Governments increased spending to stimulate the economy and ensure faster recovery of some of the other economic indicators. The series showed resilience to the shock, and the interventions by the government helped maintain or recover the relationship with other indicators. However, the interdependency between government consumption, expenditure, and investment, and import/export shifted significantly as the interventions that controlled the domestic economy could not provide control internationally. The collapse of global economic activity led to a sharp decline in trade. Emphasis on boosting domestic demand and government intervention promoted a notable decrease in aggregate expenditure, particularly in trade-intensive durable goods, and had a significant effect on imports and exports Bems et al. (2013)
The housing market bubble fueled by subprime mortgage rates before 2007 and its subsequent collapse is generally attributed to as one of the triggers for the Great Recession; Gorton (2009). However, not all housing sectors were impacted equally. During the Great Recession, single-family housing starts were severely affected. Multifamily housing experienced lower levels of rent decline and shorter recovery period until rents attained their pre-recession peaks Wang (2019) Although multifamily housing starts declined during the Great Recession, the sector’s performance in terms of rental rates and recovery time showcased its relative resilience compared to other real estate sectors. This is bourne out in the relatively fewer interdependencies between the multifamily housing and starts and other housing series that were marked as showing significant changes after the recession.
On the other hand, series such as consumer-related PCE, particularly those related to the health and education sector employment, showed resilience during the recession Barello (2014), but their interdependencies with other series seem to have been disrupted significantly by the financial crisis. This shows that while some sectors were somewhat recession proof, their interrelation with other sectors were not.
Table 2: Proportions of entries in the stationary precisions where the 95% credible interval for the associated edge-difference parameter does not include zero.
Proportion of edges that changed
National income and product accounts 0.72
Industrial Production 0.13
Employment 0.83
Unemployment and Participation 0.33
Housing 0.41
Inventories, Orders, and Sales 0.31
Prices (PCE) 0.64
Prices (CPI) 0.25
Earnings and Productivity 0.55
Interest Rates 0.43
Money and Credit 0.52
Household Balance Sheets 0.62
Exchange Rates 0.17
Stock Markets 0.57
Non-Household Balance Sheets 0.78
Figure 2: Estimated change networks. The node numbers correspond to the variable names as described in McCracken and Ng (2020).
7 Simulation